# Efficient State Preparation for Metrology and Quantum Error Correction with Global Control

### v1

@@ Abstract
We introduce a simple, experimentally realizable protocol that can prepare any specific superposition of permutationally invariant qubit states, also known as Dicke states. The protocol is comprised entirely of global rotations and globally applied non-linear phase gates — it does not require local addressability or ancilla qubits — and hence can be readily implemented in a variety of experimental platforms, including trapped-ion quantum simulators and cavity QED systems. We demonstrate the utility of our protocol by numerically preparing several states with theoretical infidelities $1-\mathcal{F}<10^{-4}$: (i) metrologically useful $N$-qubit Dicke states in $\mathcal{O}(1)$ gate steps, (ii) the $N=9$ qubit codewords of the Ruskai code with $P=4$ gate steps, and (iii) the $N=13$ qubit Gross codewords in $P=7$ gate steps. Focusing on trapped-ion platforms, we estimate that the protocol achieves fidelities $\gtrsim 95\%$ in the presence of typical experimental noise levels, thus providing a pathway to the preparation of a variety of useful highly-entangled quantum states.
The preparation of multi-partite entangled states is essential in current efforts towards quantum advantage in quantum enhanced sensing and quantum computing. In quantum enhanced sensing, multi-partite entanglement enhances the precision of the sensor from the standard quantum limit (SQL) imposed when using an uncorrelated $N$-qubit state to the Heisenberg limit [1, 2, 3, 4]. Recent impressive experimental demonstrations of the preparation of metrologically useful entangled states [5, 6, 7, 8, 9] highlight two limitations in the NISQ era: the fragility of multi-partite entanglement towards noise limits the fidelity of the state preparation, and the lack of single-site addressing restricts the variety of states that one can prepare to e.g. squeezed states.
@@
The operation of a quantum computer also relies on our ability to prepare highly entangled quantum states, both as a step for quantum algorithms, and as codewords for quantum error correction (QEC). Despite recent remarkable progress towards fault-tolerant quantum computing [10, 11, 12] there are other quantum error correcting codes which may offer advantages e.g. in terms of the types of errors they can protect against [13]. One promising approach utilizes codewords which are superpositions of states in the permutationally invariant manifold, also known as Dicke states [14, 15, 16]. Dicke states are also relevant for quantum metrology [17], including error corrected quantum sensing [18], quantum storage [19, 20] and quantum networking [21]. The importance of limiting the circuit depth in NISQ devices to realize the desired state with high enough fidelity has resulted in the recent development of many state preparation protocols which utilize ancilla qubits [22, 23], and shorter gate sequences interleaved with measurements and classical computation [24].
@@
In this paper we leverage the ability to engineer infinite range interactions in quantum simulator platforms, including trapped ions and cavity QED systems, to create arbitrary superpositions of the Dicke states with applications to quantum error correction and quantum enhanced metrology. Our protocol consists of only global operations: a global non-linear phase gate and a global rotation, and thus does not demand local addressability nor ancilla qubits. We require only $\mathcal{O}(1)$ steps to prepare metrologically useful Dicke states in tens of ions with theoretical infidelities $1-\mathcal{F}\sim 10^{-4}$. Moreover, we require $P_{\rm Ruskai}=4$ and $P_{\rm Gross}=7$ to prepare quantum error correction codewords of the $N=9$ Ruskai [15] and $N=13$ Gross [16] codes with $1-\mathcal{F}<10^{-4}$. We discuss the possible implementation in trapped ions, including expected sources of error. The limited resources required by our protocol make it immediately realizable on a broad range of experimental quantum simulation platforms.
@@
Figure 1: (a) Our protocol utilizes operations that act globally on a collection of $N$ qubits. (b) The Dicke states are permutationally invariant, which means that any qubits $i$ and $j$ can be interchanged without changing the quantum state. (c) Schematic representation of the protocol. The system is initialized in a coherent spin state (CSS) $\lvert\psi_{0}\rangle=\lvert\bar{\theta},\bar{\phi}\rangle$. Next, the one-axis twisting Hamiltonian $H_{\text{OAT}}$ generates spin squeezing. For relatively small $\phi_{k}$ the CSS is spin squeezed; for larger $\phi_{k}$ the probability distribution separates across the Bloch sphere, as is the case sketched here. Next, the state is rotated about an arbitrary axis. These two operations are repeated $P$ times, with their parameters numerically optimized to maximize the fidelity between the final state $\lvert\psi_{\text{f}}\rangle$ and a chosen target state $\lvert\psi_{\text{t}}\rangle$. Above the Bloch spheres, to highlight the protocol’s relation to amplitude amplification, we roughly sketch the $N+1$ basis state amplitudes with colouring representing the complex phase. In this perspective, $H_{\text{OAT}}$ applies a phase that depends non-linearly on the excitation number $M$, with the rotation operator then amplifying or attenuating the amplitude depending on the phase that was applied.
@@
Dicke states: We consider an ensemble of $N$ qubits. The collective spin operators are defined as $J_{\alpha}=\sum_{j=1}^{N}\sigma_{j}^{\alpha}$ where $\alpha\in\{x,y,z\}$ and $\sigma_{j}^{\alpha}$ is the Pauli operator acting on the $j$-th qubit. The collective raising and lowering operators are $J_{\pm}=J_{x}\pm iJ_{y}$. Dicke states, denoted $\lvert J,M\rangle$, are simultaneous eigenstates of the total spin operator $J^{2}=J_{x}^{2}+J_{y}^{2}+J_{z}^{2}$ and $J_{z}$. The quantum numbers $J$ and $M$ are defined by $J^{2}\lvert J,M\rangle=J(J+1)\lvert J,M\rangle$, $J_{z}\lvert J,M\rangle=M\lvert J,M\rangle$ with $J\in\{0,\ldots,N/2\}$, $M\in\{-J,\ldots,J\}$ taking (half)-integer values for even (odd) $N$. In this work we focus on the symmetric $J\equiv N/2$ Dicke states. The $N+1$ states in this manifold can be written in the single-particle $\sigma_{z}$ basis as
@@
where the normalisation factor is $\lambda=\begin{pmatrix}2J\\ J+M\end{pmatrix}$, and where the summation is over all permutations $\mathcal{P}$ of the ordering of the spins. For example, for $N=3$ the four symmetric Dicke states are |3/2,3/2=|000$\lvert$ 3/2,-3/2$\rangle$=$\lvert$ 000$\rangle$| 3 / 2 , - 3 / 2 ⟩ = | 000 ⟩, $\lvert 3/2,-1/2\rangle=1/\sqrt{3}(\lvert 001\rangle+\lvert 010\rangle+\lvert 100\rangle)$, $\lvert 3/2,1/2\rangle=1/\sqrt{3}(\lvert 011\rangle+\lvert 101\rangle+\lvert 110\rangle)$ and |3/2,3/2=|111$\lvert$ 3/2,3/2$\rangle$=$\lvert$ 111$\rangle$| 3 / 2 , 3 / 2 ⟩ = | 111 ⟩. The entanglement of Dicke states is persistent [25], which makes them attractive for experimental realization. Dicke states such as the $N$-qubit W state or the Dicke state with $N/2$ excitations are robust and demonstrate useful sensing capabilities [26, 27, 28]. There are many deterministic preparation protocols for Dicke states, including quantum circuit approaches [23, 22, 24] and global pulse schemes [29, 30, 31, 32], among others [33, 34, 35].
In addition to bare Dicke states, superpositions of Dicke states are of key interest. For example, the Greenberger–Horne–Zeilinger (GHZ) state $\lvert\text{GHZ}\rangle=\lvert N/2,-N/2\rangle+\lvert N/2,N/2\rangle$ is maximally entangled, represents a distinct class of multi-partite entanglement and exhibits phase sensing sensitivity below the SQL [36, 37]. Its preparation has been experimentally demonstrated in platforms including trapped ions [38], Rydberg atom arrays [5] and superconducting qubits [6].
Going beyond simple Dicke states, preparation of more complicated states quickly becomes challenging. This is particularly the case for the codewords of Pauli exchange error correcting codes [15] or the codewords of large spin codes [16] which are superpositions of Dicke states with highly specific coefficients.
State preparation protocol: Our protocol requires a phase gate that depends non-linearly on $M$. We use the one-axis twisting Hamiltonian $H_{\text{OAT}}=gJ_{z}^{2}$ with coupling strength $g$ [39]. The corresponding evolution operator is
$U_{\text{OAT}}(\phi)=e^{-i\phi J_{z}^{2}},$ (2)
where $\phi=gt$ with $t$ the duration for which $H_{\text{OAT}}$ is applied. We construct a gate sequence that is composed of $U_{\text{OAT}}$ interleaved with global rotations about an arbitrary axis described by $R(\theta_{k},\xi_{k},\zeta_{k})=R_{z}(\theta_{k})R_{y}(\xi_{k})R_{z}(\zeta_{k})$. The resulting unitary is
$U_{\text{tot}}=\prod_{k=1}^{P}[R(\theta_{k},\xi_{k},\zeta_{k})U_{\text{OAT}}(\phi_{k})],$ (3)
where $P$ denotes the number of steps. We note that in the large $N$ limit, after making the Holstein-Primakoff transformation [40], $H_{\text{OAT}}$ and $R$ become a Kerr nonlinearity and displacement operator, respectively, which may open a path for other experimental implementations.
We sketch a schematic overview of the protocol in Fig. 1. The protocol design is motivated by amplitude amplification: in the computational basis $U_{\text{OAT}}$ applies a phase that depends non-linearly on $M$. When the state is then rotated about a chosen axis, the amplitude of each $\lvert N/2,M\rangle$ is either amplified or attenuated, depending on the phase that was imparted by $U_{\text{OAT}}$. In this way, repeated application of $U_{\text{OAT}}$ and $R$ with optimized parameters can maximize the amplitude of a target state while minimizing the amplitude of all other states in Hilbert space. The parameters $\bm{\theta},\bm{\xi},\bm{\zeta}$ and $\bm{\phi}$ are chosen to minimize the infidelity $1-\mathcal{F}=1-|\langle\psi_{\text{t}}|\psi_{\text{f}}\rangle|^{2}$ where $\lvert\psi_{\text{t}}\rangle$ is a chosen target state, $\lvert\psi_{\text{f}}\rangle=U_{\text{tot}}\lvert\psi_{0}\rangle$ the final state and $\lvert\psi_{0}\rangle$ the initial state.
@@
Because the gate sequence Eq. (3) begins with the application of $U_{\text{OAT}}$, for the initial state we choose a global rotation acting on $\lvert N/2,-N/2\rangle$, which is a coherent spin state (CSS) $\lvert\psi_{0}\rangle=\lvert\bar{\theta},\bar{\phi}\rangle$, ie. an eigenstate of the spin component in the $(\bar{\theta},\bar{\phi})$ direction. The parameters $(\bar{\theta},\bar{\phi})$ are also numerically optimized, and thus the total number of free parameters is $34P+$ 3.
Dicke state preparation: We demonstrate numerically that the protocol Eq. (3) can prepare $N$-qubit symmetric Dicke states with high theoretical fidelity. Our numerical optimisation uses a multi-start Monte Carlo search, with local optimisation implemented using L-BFGS [41] in Julia [42] with Optim.jl and LineSearches.jl [43].
@@
Figure 2: (a) Theoretical infidelity $1-\mathcal{F}=1-|\langle\psi_{\text{t}}|\psi_{\text{f}}\rangle|^{2}$ to prepare: the $N/2$ excitation state $\lvert N/2,0\rangle$ (solid, dots) and the W state $\lvert N/2,-N/2+1\rangle$ (dashed, stars) using the protocol $U_{\text{tot}}$ Eq. (3). Both target states, even for $N=60$, can be prepared with small theoretical infidelities $1-\mathcal{F}\sim 10^{-4}$ in $P=4$ and $P=3$ gate steps respectively. (b) The su(2) Husimi Q-representation for the $P=3$ gate sequence solution that prepares $\lvert N/2,0\rangle$ with $N=60$. Starting with the spins pointing along $-x$, the interleaving of spin squeezing generated by $H_{\text{OAT}}$ and the global rotations produces a trajectory through Hilbert space that leads to the preparation of the target state with infidelity $1-\mathcal{F}\sim 2\times 10^{-3}$. The optimized parameters used for this sequence are $\phi=[0.054,4.712,3.441]$, $\zeta=[0.412,3.142,3.096]$ and $\xi=[4.474,3.142,3.115]$.
@@
Firstly, we prepare the Dicke state with $N/2$ excitations, $\lvert\psi_{\text{t}}\rangle=\lvert N/2,0\rangle$. Note that this state is metrologically useful: when the state used for sensing is $\lvert N/2,0\rangle$, the uncertainty in the measured angle $\epsilon$ of a rotation $e^{-i\epsilon J_{y}}$ saturates the Cramer-Rao bound, $(\Delta\epsilon)^{2}=2/[N(N+2)]$ and is robust to particle loss [44, 45]. We exploit the symmetry of $\lvert N/2,0\rangle$ in the Dicke manifold to reduce the number of parameters by restricting the global rotations in the protocol to $R_{z}(\theta_{k})$ and $R_{y}(\xi_{k})$, and by setting the initial CSS to be an eigenstate of $J_{x}$ such that $\langle\psi_{0}\rvert J_{x}\lvert\psi_{0}\rangle=-N/2$. In Fig. 2(a) we plot (solid lines with dots) the infidelity $1-\mathcal{F}=1-|\langle\psi_{\text{t}}|\psi_{\text{f}}\rangle|^{2}$ as a function of $N$ for $P=2$ (blue), $P=3$ (orange) and $P=4$ (green). The infidelity is remarkably robust to increasing qubit number, with $P=4$ gates sufficient to prepare the target state in tens of ions with infidelities $1-\mathcal{F}\sim 10^{-4}$. In Fig. 2(b) we plot the evolution of the $N=60$ state under the $P=3$ gate sequence solution using the su(2) Husimi Q-Representation $Q(\theta,\phi)=\langle\theta,\phi\rvert\rho\lvert\theta,\phi\rangle$ of a state $\rho$ where $\lvert\theta,\phi\rangle$ is a CSS. Clearly visible is the effect of spin squeezing, whose repeated application combined with global rotations results in the preparation of $\lvert N/2,0\rangle$.
@@
Secondly, we prepare the $N$-qubit W state, $\lvert N/2,-N/2+1\rangle$. In Fig. 2(a), we plot (dashed lines with stars) the infidelity to prepare the W state as a function of $N$ with $P=2$ (blue) and $P=3$ (orange) gate steps. Note that the superposition of all states with only one qubit not excited $\lvert N/2,N/2-1\rangle$ can be prepared using the same sequence by applying a global spin flip at the end of the sequence, and that optimizing for $\lvert N/2,N/2-1\rangle$ rather than $\lvert N/2,-N/2+1)\rangle$ does not offer any potential for improvement as the initial state is an optimized CSS. Importantly, we find that the $N=60$ W state can be prepared with small infidelities in only $P=3$ gates. Thus, our protocol can be used to prepare large $N$ Dicke states in $\mathcal{O}(1)$ gate steps.
Quantum error correction codewords: Next, we demonstrate the versatility of our protocol by preparing specific superpositions of Dicke states. We focus on the codewords of the Ruskai and Gross codes.
@@
Pauli exchange errors are two-qubit errors that arise due to interactions between qubits, which can occur for example due to dipole-dipole interactions. Physically, exchange errors are similar to bit flip errors, with the additional constraint that a qubit will flip if and only if its neighbour is different. Codes such as the $9$-qubit Shor code [46] cannot distinguish between exchange errors and phase errors. The Ruskai code [15] exploits the permutation invariance of Dicke states to correct all one qubit errors and Pauli exchange errors. The explicit codewords for $N=9$ are
$\lvert R_{0}\rangle$ $=\lvert 9/2,-9/2\rangle+\frac{1}{\sqrt{28}}\lvert 9/2,3/2\rangle,$ (4a)
$\lvert R_{1}\rangle$ $=\lvert 9/2,9/2\rangle+\frac{1}{\sqrt{28}}\lvert 9/2,-3/2\rangle,$ (4b)
where the coefficient $1/\sqrt{28}$ is required to satisfy the QEC condition [15].
@@
where $\lvert G_{1}\rangle$ follows from $\lvert G_{0}\rangle$ by a global spin flip, and where the coefficients are all determined by the error correction condition [16].

### v2

# Global Variational Quantum Circuits for Arbitrary Symmetric State Preparation

@@ Abstract
Quantum states that are symmetric under particle exchange play a crucial role in fields such as quantum metrology and quantum error correction. We use a variational circuit composed of global one-axis twisting and global rotations to efficiently prepare arbitrary symmetric states, i.e. any superposition of Dicke states. The circuit does not require local addressability or ancilla qubits and thus can be readily implemented in a variety of experimental platforms including trapped-ion quantum simulators and cavity QED systems. We provide analytic and numerical evidence that any $N$-qubit symmetric state can be prepared in $32N/$ 3 steps. We demonstrate the utility of our protocol by preparing (i) metrologically useful $N$-qubit Dicke states of up to $N=300$ qubits in $\mathcal{O}(1)$ gate steps with theoretical infidelities $1-\mathcal{F}<10^{-3}$, (ii) the $N=9$ Ruskai codewords in $P=4$ gate steps with $1-\mathcal{F}<10^{-4}$, and (iii) the $N=13$ Gross codewords in $P=7$ gate steps with $1-\mathcal{F}<10^{-4}$. Focusing on trapped-ion platforms, for the $N=9$ Ruskai and $N=13$ Gross codewords we estimate that the protocol achieves fidelities $\gtrsim 95\%$ in the presence of typical experimental noise levels, thus providing a pathway to the preparation of a wide range of useful highly-entangled quantum states.
@@
The operation of a quantum computer relies on our ability to prepare highly entangled quantum states, both as a step for quantum algorithms, and as codewords for quantum error correction (QEC). Despite recent remarkable experimental progress towards fault-tolerant quantum computing [1, 2, 3], the fragility of multi-partite entanglement to noise limits the fidelity of state preparation. Furthermore, the lack of single-site addressing in some experimental platforms has typically limited the variety of states that can be prepared to e.g. squeezed states. In light of these limitations, the choice of QEC scheme is important. Depending on its properties, certain QEC schemes may offer advantages both in terms of the types of errors that they can protect against [4], as well as their suitability of implementation to the particular hardware. One promising approach utilizes codewords which are symmetric states, i.e. superpositions of Dicke states [5, 6, 7]. The Dicke states are also relevant for quantum metrology [8], including error-corrected quantum sensing [9], quantum storage [10, 11] and quantum networking [12]. The need to overcome the intrinsic noise in NISQ devices has motivated the design of symmetric state preparation protocols that minimize the circuit depth. In particular, symmetric states have been prepared using one- and two-qubit gate circuits with depth $\mathcal{O}(N)$ [13, 14], and specific Dicke states with constant depth using measurements and classical computation [15].
@@
In this article, we demonstrate that arbitrary superpositions of Dicke states can be engineered using linear depth circuits with only global control. Specifically, we utilize a variational quantum circuit consisting of only global one-axis twisting and global rotations, and thus do not demand local addressability nor ancilla qubits. We provide analytic and numerical evidence that the number of steps to synthesize an arbitrary $N$-qubit symmetric states scales as $32N/$ 3. Focusing on applications to QEC codewords, we show that only $P=4$ and $P=7$ steps are required to prepare QEC codewords of the $N=9$ Ruskai [6] and $N=13$ Gross [7] codes with infidelity $1-\mathcal{F}<10^{-4}$, respectively. We discuss the possible implementation in trapped ions, including a detailed analysis of expected sources of error. The limited resources required by our protocol make it immediately realizable on a broad range of experimental quantum simulation platforms.
@@
Figure 1: (a) The protocol Eq. (2) utilizes operations that act globally on a collection of $N$ qubits. (b) The Dicke states are symmetric, i.e. any two qubits $i$ and $j$ can be interchanged without changing the quantum state. (c) Schematic representation of the protocol. The system is initialized in a coherent spin state (CSS) $\lvert\psi_{0}\rangle=\lvert\bar{\theta},\bar{\phi}\rangle$. Next, one-axis twisting $U_{\text{OAT}}$ generates spin squeezing. For relatively small $\phi_{k}$ the CSS is spin squeezed; for larger $\phi_{k}$ the probability distribution separates across the Bloch sphere, as is the case sketched here. Next, the state is rotated about an arbitrary axis. These two operations are repeated $P$ times, with parameters numerically optimized to minimize the infidelity $1-\mathcal{F}$. Above the Bloch spheres, to highlight the connection to amplitude amplification, we sketch the $N+1$ basis state amplitudes with colouring representing complex phase. In this perspective, $U_{\text{OAT}}$ applies a phase that depends non-linearly on the excitation number $M$, with the rotation then amplifying or attenuating the amplitude conditioned on the phase that was applied. Repeated application of $U_{\text{OAT}}$ and $R$ with optimized parameters can therefore maximize the amplitude of a target state.
@@
Dicke states: We consider an ensemble of $N$ qubits. The collective spin operators are defined as $J_{\alpha}=(1/2)\sum_{j=1}^{N}\sigma_{j}^{\alpha}$ where $\alpha\in\{x,y,z\}$ and $\sigma_{j}^{\alpha}$ is the Pauli operator acting on the $j$-th qubit. Dicke states, denoted $\lvert J,M\rangle$, are simultaneous eigenstates of the total spin operator $J^{2}=J_{x}^{2}+J_{y}^{2}+J_{z}^{2}$ and $J_{z}$. The quantum numbers $J$ and $M$ are defined by $J^{2}\lvert J,M\rangle=J(J+1)\lvert J,M\rangle$, $J_{z}\lvert J,M\rangle=M\lvert J,M\rangle$ with $J\in\{0,\ldots,N/2\}$, and $M\in\{-J,\ldots,J\}$. We focus on the $N+1$ symmetric $J\equiv N/2$ Dicke states, which in the single-particle $\sigma_{z}$ basis are
@@
where the normalisation factor is $\lambda=\begin{pmatrix}2J\\ J+M\end{pmatrix}$, and where the summation is over all permutations $\mathcal{P}$ of the ordering of the spins  [^1] For example, the four $N=3$ symmetric Dicke states are |3/2,3/2=|000$\lvert$ 3/2,-3/2$\rangle$=$\lvert$ 000$\rangle$| 3 / 2 , - 3 / 2 ⟩ = | 000 ⟩, $\lvert 3/2,-1/2\rangle=1/\sqrt{3}(\lvert 001\rangle+\lvert 010\rangle+\lvert 100\rangle)$, $\lvert 3/2,1/2\rangle=1/\sqrt{3}(\lvert 011\rangle+\lvert 101\rangle+\lvert 110\rangle)$ and |3/2,3/2=|111$\lvert$ 3/2,3/2$\rangle$=$\lvert$ 111$\rangle$| 3 / 2 , 3 / 2 ⟩ = | 111 ⟩. Dicke states feature entanglement that is persistent [17] and robust to particle loss [18, 19, 20] which makes them attractive for experimental realization, while also demonstrating useful sensing capabilities [21, 22]. Specifically, the $N$-qubit W-state $\lvert N/2,-N/2+1\rangle$ is maximally pairwise entangled [23, 24], while the Dicke state with $N/2$ excitations $\lvert N/2,0\rangle$ saturates the quantum Cramér-Rao bound when sensing global rotations about the x-y axis [25]. There are many existing deterministic preparation protocols for Dicke states, including quantum circuit approaches that use one- and two-qubit gates [14, 13, 15], global pulse schemes [25, 26, 27, 28], and others [29, 30, 31].
In addition to bare Dicke states, superpositions of Dicke states are of key interest. For example, the Greenberger–Horne–Zeilinger (GHZ) state $\lvert\text{GHZ}\rangle=\lvert N/2,-N/2\rangle+\lvert N/2,N/2\rangle$ is maximally entangled, represents a distinct class of multi-partite entanglement, exhibits phase sensing sensitivity below the standard quantum limit (SQL) [32, 33] and has been prepared in many experimental platforms [34, 35, 36]. More generally, superpositions over more of the Dicke states form the logical basis states of various QEC codes including Pauli exchange error correcting codes [6] and large spin codes [7]. The weighting of individual Dicke states in each of these logical basis states are chosen to satisfy the Knill-Laflamme QEC conditions [37]. Existing protocols to prepare such specific, highly entangled superpositions of Dicke states typically require single-site addressability [38] or a large detuning limit [39]. Below, we alleviate these constraints and propose to prepare Dicke state superpositions using a global variational quantum circuit.
State preparation protocol: Global variational quantum circuits have been previously proposed [40, 41, 42, 43] and demonstrated experimentally [44] for the preparation of squeezed quantum states for quantum metrology. Here, instead of optimizing the variational parameters to maximize a property of the target state (such as the squeezing amplitude or phase sensitivity), we employ such circuits to prepare target symmetric quantum states by minimizing the infidelity $1-\mathcal{F}=1-|\langle\psi_{\text{t}}|\psi_{\text{f}}\rangle|^{2}$, where $\lvert\psi_{\text{t}}\rangle$ is a chosen target state, $\lvert\psi_{\text{f}}\rangle=U_{\text{tot}}\lvert\psi_{0}\rangle$ the final state and $\lvert\psi_{0}\rangle$ the initial state. As sketched in Fig. 1, the circuit of $P$ layers is
$U_{\text{tot}}=\prod_{k=1}^{P}[R(\theta_{k},\xi_{k})U_{\text{OAT}}(\phi_{k})],$ (2)
and consists of: (i) global rotations $R(\theta_{k},\xi_{k})=R_{z}(\theta_{k})R_{y}(\xi_{k})$, and (ii) global one-axis twisting $U_{\text{OAT}}(\phi)=e^{i\phi J_{z}^{2}}$, where $\phi=gt$ with $g$ the coupling strength and $t$ the evolution time [45]. Because $R_{z}$ commutes with $U_{\text{OAT}}$, for $P\geq 2$ the global rotation $R$ is about an arbitrary axis. Although the fidelity can be efficiently computed classically because the symmetric space dimension is $N+1$, in the Supplemental Material we discuss how an upper bound on the infidelity can be efficiently computed on quantum devices for a hybrid classical-quantum algorithm approach, which may be a preferred option to mitigate certain hardware errors [46].
The capability of the protocol to prepare superpositions of Dicke states can be understood using amplitude amplification, as sketched in Fig. 1. To emphasize the versatility of the variational circuit, we note that $U_{\text{OAT}}$ and $R$ provide universal control over the symmetric subspace [47, 43]. Although a general explicit expression for the optimal parameters for arbitrary symmetric state synthesis is a complicated task, a simple consistency argument for the minimum number of gates $P^{*}$ to prepare a general symmetric state can be made as follows: accounting for the coefficients and normalization of the $N+1$ basis states, a general state in the symmetric $J=N/2$ subspace is described by $2N$ real parameters. As each protocol step is characterized by three free parameters $\theta,\xi,\phi$, the preparation of an arbitrary symmetric state requires $P^{*}=2N/3$ steps.
We numerically verify the scaling relation for $P^{*}$ by generating $N_{\text{Haar}}=200$ Haar random states as target states and optimizing the protocol’s variational parameters to minimize $1-\mathcal{F}$. Obtaining (and verifying) a global minima is challenging, as in general the optimization landscape can contain many local minima. However, if the system is controllable [48], then every local minima is a global minima [49]. As such, performing only local optimization of the parameters we expect a sharp transition in target state infidelity at some $P\geq P^{*}$ as the system becomes controllable.
Figure 2: Infidelity $1-\mathcal{F}=1-|\langle\psi_{\rm t}|\psi_{\rm f}\rangle|^{2}$ vs. circuit depth $P$ for various $N$. Each data point is a Haar random target state; we draw $N_{\text{Haar}}=200$ targets for each $N$. The variational parameters are classically optimized using a local optimization algorithm. As the system becomes controllable, all local minima become global minima, which corresponds for each $N$ to the sharp decrease in infidelity as $P$ increases (marked with vertical dashed lines). Inset: location of the transition as a function of $N$ (black points) for up to $N=250$, demonstrating excellent agreement with the scaling prediction $P^{*}=2N/3$ (grey dashed).
In Fig. 2 we plot the infidelity $1-\mathcal{F}$ as a function of $P$ for various $N$ up to $N=50$. We perform local gradient-based optimisation using BFGS [50] in Julia [51] with Optim.jl and LineSearches.jl [52], and terminate the optimization once the infidelity passes below a threshold $\epsilon=10^{-12}$. As expected, as $P$ increases the infidelity decreases, with a relatively sharp transition as the system becomes controllable. In the inset of Fig. 2 we plot the location of the transition as a function of $N$, showing excellent agreement with the scaling obtained from counting, $P^{*}=2N/3$. As shown in the inset we further verify this prediction by optimizing the parameters for $N_{\rm Haar}=200$ Haar random states using $P=\lceil 2N/3\rceil+c$ gates with a small constant offset $c$  [^2] An offset $c$ in the number of gates $P=\lceil 2N/3\rceil+c$ is needed to reach fidelities below the numerical termination threshold $\epsilon=10^{-12}$. For $N=10$ to $N=150$ the offset is either $c=2$ or $c=3$, depending on the ceil function. For $N=200$ and $N=250$ the offset is $c=4$, to give the numerical optimization more overhead to reach the termination threshold. Specifically, for $N=[100,150,200,250$ we use $P=[70,103,138,171]$ gates. for $N=100$, $N=150$, $N=200$ and $N=250$, verifying that all states can be prepared with infidelities below the optimization termination threshold $\epsilon=10^{-12}$. Therefore the protocol Eq. (2) is an efficient preparation protocol for synthesizing arbitrary symmetric states. To the best of our knowledge this improves upon previous global control protocols, which required $N$ steps only in a large detuning limit [39] or using excited Rydberg states [54], and $\mathcal{O}(N^{4})$ steps for arbitrary unitary synthesis [43].
@@
Dicke state preparation: Next, we demonstrate the preparation of specific states, beginning with bare Dicke states. Because the gate sequence Eq. (2) starts with the application of $U_{\text{OAT}}$, for the initial state we choose a global rotation acting on $\lvert N/2,-N/2\rangle$, which is a coherent spin state (CSS) $\lvert\psi_{0}\rangle=\lvert\bar{\theta},\bar{\phi}\rangle$, i.e. an eigenstate of the spin component in the $(\bar{\theta},\bar{\phi})$ direction. The parameters $(\bar{\theta},\bar{\phi})$ are also numerically optimized, and thus the total number of free parameters is $23P+$ 2. Both here and in the next section, the numerical optimisation uses a multi-start Monte Carlo search, with local optimization using L-BFGS [55]
@@
Figure 3: (a) Theoretical infidelity $1-\mathcal{F}=1-|\langle\psi_{\text{t}}|\psi_{\text{f}}\rangle|^{2}$ to prepare: the the W state $\lvert N/2,-N/2+1\rangle$ (solid blue) and $N/2$ excitation state $\lvert N/2,0\rangle$ (dashed orange) using the protocol $U_{\text{tot}}$ Eq. (2) for $P=2$ (dots), $P=3$ (diamonds) and $P=4$ (squares). The lines show a power law fit, with excellent agreement. Both target states can be prepared with small theoretical infidelities $1-\mathcal{F}<10^{-3}$ in $P=3$ and $P=4$ gate steps respectively in systems as large as $N=300$ qubits. (b) The su(2) Husimi Q-representation for the $P=3$ solution that prepares $\lvert N/2,0\rangle$ with $N=300$. The optimized initial coherent spin state starts along $x$, with the subsequent interleaving of one-axis twisting and global rotations leading to the preparation of $\lvert\psi_{t}\rangle$ with $1-\mathcal{F}\sim 10^{-2}$.
@@
Firstly, we prepare the Dicke state with $N/2$ excitations, $\lvert\psi_{\text{t}}\rangle=\lvert N/2,0\rangle$. This state is metrologically useful: when the state used for sensing is $\lvert N/2,0\rangle$, the uncertainty in the measured angle $\epsilon$ of a rotation $e^{-i\epsilon J_{y}}$ saturates the Cramer-Rao bound, $(\Delta\epsilon)^{2}=2/[N(N+2)]$ and features entanglement that is robust to particle loss [56, 57, 18, 19, 20]. In Fig. 3(a) we plot (dashed orange) $1-\mathcal{F}$ as a function of $N$ for $P=2$ (circles), $P=3$ (diamonds) and $P=4$ (squares). The infidelity is remarkably robust to increasing qubit number, with $P=4$ gates sufficient to prepare the target state in up to $N=300$ qubits with $1-\mathcal{F}\lesssim 10^{-2}$. The dashed lines show a power law fit with excellent agreement, emphasizing the slow increase of infidelity with $N$ for $N\gtrsim 50$. In Fig. 3(b) we plot the evolution of the $N=300$ state under the $P=3$ gate sequence solution using the su(2) Husimi Q-Representation $Q(\theta,\phi)=\langle\theta,\phi\rvert\rho\lvert\theta,\phi\rangle$ of a state $\rho$ where $\lvert\theta,\phi\rangle$ is a CSS. Clearly visible is the effect of spin squeezing, whose repeated application combined with global rotations results in the preparation of $\lvert N/2,0\rangle$.
@@
Secondly, we prepare the $N$-qubit W state, $\lvert N/2,-N/2+1\rangle$. In Fig. 3(a), we plot (solid blue) the infidelity to prepare the W-state with $P=2$ (circles) and $P=3$ (diamonds)  [^3] The superposition of all states with only one qubit not excited $\lvert N/2,N/2-1\rangle$ can be prepared using the same sequence by applying a global spin flip at the end of the sequence, and that optimizing for $\lvert N/2,N/2-1\rangle$ rather than $\lvert N/2,-N/2+1\rangle$ does not offer any potential for improvement as the initial state is an optimized CSS.. We find that the W state can be prepared in systems of up to $N=300$ qubits with small infidelities $1-\mathcal{F}<10^{-4}$ in only $P=3$. Thus, our protocol can be used to prepare these two Dicke states in up to $N=300$ qubits in $\mathcal{O}(1)$ gate steps.
@@
Quantum error correction codewords: Next, we prepare specific superpositions of Dicke states for QEC. Pauli exchange errors are two-qubit errors that arise due to interactions between qubits, for example due to dipole-dipole interactions. Physically, exchange errors are similar to bit flip errors, with the additional constraint that a qubit will flip if and only if its neighbour is different. Codes such as the $9$-qubit Shor code [59] cannot distinguish between exchange errors and phase errors. The Ruskai code [6] exploits the permutation invariance of Dicke states to correct all one qubit errors and Pauli exchange errors. The explicit codewords for $N=9$ are
$\lvert R_{0}\rangle$ $=\lvert 9/2,-9/2\rangle/2+\sqrt{3/4}\lvert 9/2,3/2\rangle,$ (3a)
$\lvert R_{1}\rangle$ $=\lvert 9/2,9/2\rangle/2+\sqrt{3/4}\lvert 9/2,-3/2\rangle,$ (3b)
@@
where $\lvert G_{1}\rangle$ follows from $\lvert G_{0}\rangle$ by a global spin flip.